# LSAQ: Layer-Specific Adaptive Quantization for Large Language Model Deployment

### v1

@@ Abstract
As large language models (LLMs) demonstrate exceptional performance across various domains, the deployment of these models on edge devices has emerged as a new trend. Quantization techniques, which reduce the size and memory footprint of LLMs, are effective for enabling deployment on resource-constrained edge devices. However, existing one-size-fits-all quantization methods often fail to dynamically adjust the memory consumption of LLMs based on specific hardware characteristics and usage scenarios. To address this limitation, we propose LSAQ (Layer-Specific Adaptive Quantization), a system for adaptive quantization and dynamic deployment of LLMs based on layer importance. LSAQ evaluates layer importance by constructing top-$k$ token sets from the inputs and outputs of each layer and calculating their Jaccard coefficient. Using this evaluation, the system adaptively adjusts quantization strategies in real time according to the resource availability of edge devices, assigning different precision levels to layers of varying importance. This approach significantly reduces the storage requirements of LLMs while maintaining model performance, enabling efficient deployment across diverse hardware platforms and usage scenarios.
@@
large language models, quantization, layer importance, jaccard coefficient, edge devices
@@ I Introduction
With the rapid advancement of large language models (LLMs), they have demonstrated exceptional capabilities across a variety of tasks. Consequently, the use of LLMs has proliferated across multiple domains, including natural language processing [1], code generation [2, 3], and even in finance [4] and education [5] sectors. To harness LLMs for more intelligent and personalized services [6] while ensuring data security and privacy, the deployment of LLMs on edge devices has emerged as a new trend. However, the massive parameter size of these models poses significant challenges for efficient deployment on edge devices. For example, a model with tens of billions of parameters, such as Llama-2-13B [7], requires approximately 25 GB of memory just to load its weights. In contrast, a high-performance consumer-grade GPU such as the NVIDIA GeForce RTX 4090 has only 24 GB of memory, rendering it incapable of deploying Llama-2-13B directly. To effectively address this challenge, researchers are exploring model compression techniques such as quantization [8, 9], pruning [10, 11, 12], distillation [13, 14], and low-rank factorization [15] to optimize the storage and computational requirements of LLMs.
@@
In the field of model compression, model quantization plays a pivotal role. Its core mechanism involves converting high-precision floating-point weights or activation values in a model into lower-precision numerical representations. This transformation significantly reduces the size of the model storage and minimizes memory consumption during inference. Post-training quantization (PTQ), a widely adopted quantization technique, is particularly valuable for deploying models on resource-constrained devices, such as mobile devices and embedded systems. Quantization methods such as LLM.int8() [9], GPTQ [8], and Omniquant [16] fall under the PTQ category.
@@
However, many prevalent PTQ methods uniformly quantize across all layers of LLMs, neglecting the varying significance of these layers. This practice disregards evidence from existing studies [21] that indicate substantial redundancy in numerous layers within LLMs. This one-size-fits-all quantization approach also limits the ability to dynamically adjust the memory consumption of LLMs based on specific hardware characteristics and task requirements. The concept of Layer-Wise Quantization [17] offers a potential solution to this problem by quantifying each layer based on the cosine similarity between the input and output of the layer, or the distribution of weight outliers, as indicators of layer importance. Nevertheless, these methods have limitations: neither directly reflects the degree of semantic change, and the latter struggles to effectively distinguish layers with relatively uniform weight distributions. Therefore, it is crucial to explore more refined quantization strategies to enable differentiated treatment of LLM layers. Additionally, techniques capable of dynamically adjusting the quantization precision based on practical application scenarios should be developed to further optimize the storage and inference efficiency while maintaining model performance.
Figure 1: Framework of LSAQ system
@@
To that end, we propose LSAQ, a novel Layer-Specific Adaptive Quantization framework tailored for the deployment of LLMs, as illustrated in Figure 1. LSAQ introduces a fine-grained layer importance evaluation mechanism by constructing top-$k$ token sets corresponding to the inputs and outputs of each layer in LLMs. It calculates the Jaccard coefficient between these two sets as an indicator of layer importance. A higher Jaccard coefficient implies greater similarity between the input and output token sets, suggesting that the layer has not achieved significant semantic transformation when processing the input. Consequently, such layers are deemed less important. Based on the resource constraints of edge devices, the system adaptively adjusts the quantization strategy, assigning relatively higher precision to layers with greater importance and lower precision to less important layers. Additionally, when device resources allow, the system maximizes the number of high-precision layers to preserve model performance. Finally, the quantized LLMs are deployed to edge devices to handle various tasks. This intelligent adaptive quantization approach significantly reduces the storage requirements of LLMs while maintaining performance. It also enables efficient deployment of LLMs across diverse hardware platforms and various usage scenarios on the same platform. The contributions of this work can be summarized as follows:
  • A novel layer importance assessment method for LLMs is proposed, which effectively captures the relationships between semantic information across layers. This method helps to reveal the features and aspects that the model focuses on at different levels, enabling a more precise evaluation of layer importance.
  • An approach has been designed for the layer-specific quantization and dynamic deployment of LLMs based on layer importance and edge device resources, which allows for the deployment of the quantized model while striving to maintain the model’s performance.
  • LSAQ surpasses existing quantization methods of the same granularity in accuracy on 87.5% of zero-shot tasks and demonstrates superior average accuracy across all tasks. Additionally, in 90% of quantization scenarios, LSAQ outperforms comparable methods in terms of perplexity. These experimental results validate the effectiveness of LSAQ in preserving essential information while achieving efficient quantization.
@@ II-A LLMs Quantization
Model quantization is a critical technique in the field of deep learning. It achieves significant model size reduction by converting model parameters (e.g., weights and activation values) from high-precision formats (typically 32-bit or 16-bit floating-point numbers) to lower-precision formats (e.g., 8-bit or 4-bit integers). This makes it feasible to deploy LLMs on resource-constrained edge devices.
There are two primary approaches to LLMs quantization: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT incorporates quantization operations during the model training phase. It uses high-precision parameters for backpropagation and low-precision parameters for forward propagation, enabling the model to adapt to the changes introduced by quantization. In contrast, PTQ is applied after model training is complete. It reduces the representation precision of weights and activation values to shrink the model size. Compared to QAT, PTQ is more suitable for scenarios requiring rapid deployment in resource-constrained environments.
In recent research, PTQ has made significant advancements in reducing memory usage and improving computational efficiency for LLMs. The ZeroQuant [18] technique supports INT8 symmetric quantization and allows different quantization levels to be set for the final model, adapting to diverse model and task requirements. The GPTQ technique determines quantization parameters using calibration data, achieving the goal of reducing model size while maintaining accuracy. LLM.int8() improves quantization efficiency by handling activation outliers through mixed-precision decomposition. The SmoothQuant [19] technique introduces a mathematically equivalent per-channel scaling transformation to smooth activations and their corresponding weights across different channels, making the model more quantization-friendly. AWQ [20] optimizes the LLM quantization process by preserving significant weights and protecting prominent ones.
The development of these techniques has collectively facilitated the deployment and application of large language models in resource-constrained environments, offering new insights and methods for optimizing and deploying deep learning models.
Figure 2: The process of constructing top-$k$ token sets
@@ II-B Layer Importance Assessment for LLMs
In the field of research on LLMs, prior studies like shortGPT [21] have revealed that certain layers within these models may perform similar or redundant operations. These layers exhibit a high degree of functional overlap in the information processing pipeline, leading to redundancy in their contributions to the model’s final output. By thoroughly analyzing the contribution of each layer to the model, we can determine the relative importance of individual layers. This analysis not only helps identify layers with minimal impact on model performance but also provides a theoretical foundation for model compression.
When exploring the importance of layers in LLMs, Men et al. [21] and Dumitru et al. [22] have proposed metrics such as Block Influence (BI) and Layer Redundancy (LR) to define layer importance. Both studies share a core mechanism based on calculating the cosine similarity between the input and output of each layer. Cosine similarity measures the angular similarity between two non-zero vectors; values closer to 1 indicate higher similarity, implying that the operations performed by the layer are more redundant or overlapping. In addition to using cosine similarity as a metric, Dumitru et al. [17] also estimates layer importance by analyzing the number of weights significantly exceeding the average value within a layer. If a layer contains many weights with absolute values much larger than the average weight of that layer, it is likely to have a greater impact on the model’s output.
By evaluating layer importance, the contribution of each layer to the overall functionality of the model can be quantitatively assessed. This approach provides a clear framework for analyzing and comparing the roles of different layers, facilitating the identification and optimization of key layers. As a result, it helps enhance the model’s efficiency and performance.
@@ III Method
This section introduces LSAQ, a system designed for layer-specific adaptive quantization and dynamic deployment of LLMs on edge devices equipped with consumer-grade GPUs. As shown in Figure 1, the system is composed of both offline and online components. Due to variations in layer importance across different LLMs, as well as differences in GPU resources available on various edge devices or even on the same device under different conditions, the offline component must determine the most suitable quantization strategy based on these factors. During the online phase, the selected quantization strategy is applied to quantize the LLM to the corresponding bit widths. The quantized LLM is then deployed onto the target device, ensuring that the model can operate efficiently in resource-constrained environments while maintaining a high level of accuracy.
@@ III-A Layer Importance Detection Module
The Layer Importance Detection (LID) module identifies the key layers that have the greatest impact on model performance by analyzing the structure and behavior of the model. This ensures that essential functions of the model are preserved during quantization.
To assess layer importance, we utilize intrinsic model features as metrics. We analyze each layer of the LLMs, extracting the hidden states at the last time step of each layer, denoted as $X_{i,in}$ and $X_{i,out}$ These hidden states capture the encoded information for the last token in the input sequence at that particular layer. Next, we multiplied these hidden states by the transpose of the embedding matrix $W_{E}^{\top}$, which maps discrete vocabulary words to a continuous vector space. This multiplication operation produces a projection of each word across the vocabulary.
Following this, we ranked the projection results to determine the indices of each word, sorted from highest to lowest based on projection values. By selecting the top-$k$ indices from the sorted results, we identified the top-$k$ most probable words for the current layer’s input and output. Finally, we converted these indices back into actual words, forming top-$k$ token sets that represent the input and output of the layer. Figure 2 illustrates the detailed process of constructing the top-$k$ token sets.
@@
After obtaining the two sets $C_{i,in}$ and $C_{i,out}$, we calculate the Jaccard similarity between the two sets as a metric for assessing the importance of the layer. Jaccard similarity [23] is widely used in various fields, including machine learning and text mining, to evaluate the similarity between sets of samples or features. This similarity is calculated based on the ratio of the size of the intersection of the two sets to the size of their union. This metric is simple yet effective, providing a clear measure of the compositional similarity between the two top-$k$ token sets.
@@
According to Equation (3), we employ the Jaccard similarity coefficient to quantify the semantic transformation degree of each layer in LLMs. Specifically, we calculate the Jaccard similarity between two top-$k$ token sets derived from the input and output of each layer and transform this value by inversion and addition to obtain the importance metric $I_{i}$ for each layer. As the Jaccard similarity coefficient increases, the importance metric $I_{i}$ decreases correspondingly.
This observation indicates that a higher similarity between the two top-$k$ token sets implies that the layer may have performed less significant semantic transformation on the input information. Consequently, it can be inferred that the layer’s contribution to semantic transformation is relatively low, leading to a lower assessment of its importance. Through Figure 3, we can visually observe the contribution of different layers to the overall performance of the Llama-2 (7B/13B) [7] and Llama-3-8B [24] models within this evaluation framework.
Figure 3: Importance of LLMs Layer
Compared to directly using cosine similarity, our approach offers a finer grained analysis by directly mapping hidden states to the vocabulary. This method captures the relationship between each layer and semantic information more effectively, helping to reveal the features and aspects each layer focuses on. As a result, it enables a more accurate evaluation of layer importance.

### v2

@@ Abstract
As Large Language Models (LLMs) demonstrate exceptional performance across various domains, deploying LLMs on edge devices has emerged as a new trend. Quantization techniques, which reduce the size and memory requirements of LLMs, are effective for deploying LLMs on resource-limited edge devices. However, existing one-size-fits-all quantization methods often fail to dynamically adjust the memory requirements of LLMs, limiting their applications to practical edge devices with various computation resources. To tackle this issue, we propose Layer-Specific Adaptive Quantization (LSAQ), a system for adaptive quantization and dynamic deployment of LLMs based on layer importance. Specifically, LSAQ evaluates the importance of LLMs’ neural layers by constructing top-$k$ token sets from the inputs and outputs of each layer and calculating their Jaccard similarity. Based on layer importance, our system adaptively adjusts quantization strategies in real time according to the computation resource of edge devices, which applies higher quantization precision to layers with higher importance, and vice versa. Experimental results show that LSAQ consistently outperforms the selected quantization baselines in terms of perplexity and zero-shot tasks. Additionally, it can devise appropriate quantization schemes for different usage scenarios to facilitate the deployment of LLMs.
@@
large language models, quantization, layer importance, Jaccard similarity, edge devices
@@ I Introduction
With the rapid advancement of Large Language Models (LLMs), they have demonstrated exceptional capabilities across a variety of tasks. Taking advantage of these capabilities, LLMs have been widely studied in multiple research domains like natural language processing [1], code generation [2, 3], finance [4] and education [5]. To harness LLMs for more intelligent and personalized services [6] while ensuring data security and privacy, deploying LLMs on edge devices has emerged as a new trend. However, LLMs’ massive parameters pose significant challenges for efficiently deploying them on resource-limited edge devices. For example, Llama-2-13B [7] requires about 25GB GPU memory when loading its weights. In contrast, a high-performance consumer-grade GPU such as the RTX 4090 has only 24 GB of memory, rendering it incapable of deploying Llama-2-13B directly. To address this challenge, researchers explore LLM compression techniques such as quantization [8, 9, 10], pruning [11, 12, 13], distillation [14, 15], and low-rank factorization [16] to reduce GPU memory requirements when deploying LLMs.
@@
In the field of LLM compression, quantization plays a pivotal role. Its core mechanism involves converting high-precision floating-point weights or activation values included in an LLM into lower-precision numerical representations. This transformation significantly reduces the memory requirements of LLM deployment. Post-Training Quantization (PTQ), a widely studied quantization technique, is particularly valuable for deploying LLMs on resource-limited devices, such as personal computers. A number of PTQ methods have achieved great success, such as LLM.int8() [9], GPTQ [8], and Omniquant [17].
@@
However, the majority of existing PTQ methods uniformly quantize across all neural layers of LLMs, neglecting the fact that these layers are of quite different importance [18]. For example, substantial redundancy exists in numerous LLMs’ neural layers. This one-size-fits-all quantization approach also limits the ability to dynamically adjust the memory requirements when deploying LLMs to various edge devices with quite different computation resources. Layer-Wise Quantization (LWQ) [19] offers a potential solution to this problem by quantifying each layer based on layer importance, which is calculated by the cosine similarity between the input and output of the layer, or the distribution of weight outliers. Despite the success of LWQ, it exposes nonnegligible limitations, i.e., it cannot reflect the degree of semantic change directly and struggles to effectively distinguish layers with relatively uniform weight distributions. In light of this, it is crucial to explore more refined quantization strategies to enable differentiated treatment of LLM layers. Additionally, techniques should be developed to dynamically adjust the quantization precision. This should be based on practical application scenarios to further optimize storage and inference efficiency, while maintaining LLM performance.
@@
To tackle the above issues, we propose LSAQ, a novel Layer-Specific Adaptive Quantization system tailored for the deployment of LLMs, as illustrated in Figure 1. LSAQ introduces a fine-grained layer importance evaluation mechanism by constructing top-$k$ token sets corresponding to the inputs and outputs of each layer in LLMs. We propose to calculate the Jaccard similarity between these two sets as the indicator of layer importance. A higher Jaccard similarity implies greater similarity between the input and output token sets, indicating that the layer fails to significantly encode the semantics when processing the input. Motivated by the above fact, we regard such layers as less important. Next, our system adaptively adjusts the quantization strategy according to the available computation resource of an edge device, applying higher quantization precision to layers with greater importance and lower precision to less important layers. Additionally, our system is capable of maximizing the number of high-precision layers to preserve LLM performance, which enables the full usage of computation resources. Finally, the quantized LLMs are deployed to edge devices to handle various tasks.
Figure 1: The framework of LSAQ. It is composed of offline and online parts. In the offline part, the importance of each layer of the LLM is first obtained, and the available GPU resources at the current moment are detected simultaneously. Based on this, a quantization strategy is meticulously formulated. Subsequently, this quantization strategy is transmitted to the online part, where the model is quantized according to this strategy.
We conducted experiments on Llama models with parameter sizes ranging from 7B to 13B, and the results show that LSAQ outperformed existing same-granularity quantization method in accuracy across 75% of the selected zero-shot tasks, and demonstrated superiority in average accuracy across all tasks. Additionally, in 90% of the cases, LSAQ performed better than the same-granularity quantization method on the perplexity metric based on the WikiText2 [20] dataset. Furthermore, we carried out deployment tests on various devices and found that LSAQ was able to select the optimal deployment scheme for the deployment of LLMs.
The contributions of this work can be summarized as follows: (1) As far as we know, we are the first to use the Jaccard similarity to calculate the importance of LLMs’ neural layers. Experimental results reveal its superiority to previous methods like cosine similarity (See §$\S$§IV-B); (2) Based on layer importance, we propose a layer-specific quantization strategy, enabling us to dynamically deploy LLMs to various edge devices with different computation resources, meanwhile maintaining the LLMs’ capabilities; (3) Compared to methods with the same quantization granularity, LLMs using LSAQ quantization to achieve better performance in terms of accuracy on zero-shot tasks and perplexity based on WikiText2.
@@ II-A LLMs Quantization
Quantization is a critical technique in the field of deep learning. It significantly reduces the model size by converting model parameters from high-precision formats (e.g., 32-bit or 16-bit floating-point numbers) to lower-precision formats (e.g., 8-bit or 4-bit integers), making it feasible to deploy LLMs on resource-limited edge devices.
There are two typical approaches for LLM quantization: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT incorporates quantization operations during LLMs’ training phase. It uses high-precision parameters for backpropagation and low-precision parameters for forward propagation, enabling LLMs to adapt to the changes introduced by quantization. In contrast, PTQ is applied after finishing model training. It reduces the representation precision of weights and activation values to reduce LLMs’ size. Compared to QAT, PTQ is more suitable for scenarios requiring rapid LLM deployment on resource-limited devices.
Recently, PTQ has made significant advancements in reducing memory usage and improving computational efficiency for LLMs. The ZeroQuant [21] technique supports INT8 symmetric quantization and allows different quantization levels to be set for the final model, adapting to diverse model and task requirements. The GPTQ [8] technique determines quantization parameters using calibration data, achieving the goal of reducing model size while maintaining accuracy. LLM.int8() [9] improves quantization efficiency by handling activation outliers through mixed-precision decomposition. The SmoothQuant [22] technique introduces a mathematically equivalent per-channel scaling transformation to smooth activations and their corresponding weights across different channels, making the model more quantization-friendly. AWQ [23] optimizes the LLM quantization process by preserving significant weights and protecting prominent ones.
The development of these techniques has collectively facilitated the deployment and application of large language models in resource-limited environments, offering new insights and methods for optimizing and deploying LLMs.
Figure 2: The process of constructing top-$k$ token sets.
@@ II-B Layer Importance Assessment for LLMs
In the field of research on LLMs, prior studies like shortGPT [18] have revealed that certain layers within these models may perform similar or redundant operations. These layers exhibit a high degree of functional overlap in the information processing pipeline, leading to redundancy in their contributions to the LLM’s final output. By thoroughly analyzing the contribution of each layer to the LLM, we can determine the relative importance of individual layers. This analysis not only helps identify layers with minimal impact on LLM performance but also provides a theoretical foundation for LLM compression.
When exploring the importance of layers in LLMs, Men et al. [18] and Dumitru et al. [24] have proposed metrics such as block influence and layer redundancy to define layer importance. Both studies share a core mechanism based on calculating the cosine similarity between the input and output of each layer. Cosine similarity measures the angular similarity between two non-zero vectors, where values closing to 1 indicate higher similarity, implying that the operations performed by the layer are more redundant or overlapping. In addition to using cosine similarity as a metric, Dumitru et al. [19] also estimate the layer importance by analyzing the number of weights significantly exceeding the average value within a layer. If a layer contains many weights with absolute values much larger than the average weight of that layer, it is likely to have a greater impact on the LLM’s output.
By evaluating layer importance, the contribution of each layer to the overall functionality of the LLM can be quantitatively assessed. This approach provides a clear framework for analyzing and comparing the roles of different layers, facilitating the identification and optimization of key layers. As a result, it helps enhance the LLM’s efficiency and performance.
@@ III Method
Input: $\mathcal{L}$: List of layer importance (in order of increasing importance), $\mathcal{M}_{ava.}$: Memory of currently available GPU
Output: $\mathcal{L}_{fp16}$, $\mathcal{L}_{int8}$, $\mathcal{L}_{int4}$: Lists of layer’s label corresponding to different precisions
1
[…]
3
4if $\mathcal{M}_{ava.}$geq Memory required for FP16 precision then
5       $\mathcal{L}_{fp16}$ = $\mathcal{L}$;
6
7else if $\mathcal{M}_{ava.}$geq Memory required for INT8 precision then
8       $\mathcal{L}_{int8}$ = $\mathcal{L}$;
9
10else if $\mathcal{M}_{ava.}$geq Memory required for INT4 precision then
11       set $len$ to the number of layers in LLM;
12       set $\mathcal{M}_{int8}$ to the difference between the $\mathcal{M}_{ava.}$ and memory required for INT4 precision;
13       set $\mathcal{S}_{8to4}$ to the memory saved by quantizing a single layer from INT8 to INT4 precision;
14       $\mathcal{N}_{int4}$ = $len$ $\mathcal{M}_{int8}$ / $\mathcal{S}_{8to4}$;
15       $\mathcal{L}_{int4}$ = $\mathcal{L}[\ :\mathcal{N}_{int4}]$;
16       $\mathcal{L}_{int8}$ = $\mathcal{L}[\mathcal{N}_{int4}:$$len$]]]];
17
18else
19       pause quantization process until other processes release memory;
20
21 end if
22
23return $\mathcal{L}_{fp16}$, $\mathcal{L}_{int8}$, $\mathcal{L}_{int4}$;
24
Algorithm 1 Quantization Precision Allocation
This section introduces LSAQ, a system designed for Layer-Specific Adaptive Quantization and dynamic deployment of LLMs on edge devices equipped with consumer-level GPUs. As shown in Figure 1, our system is composed of both offline and online components. Due to variations in layer importance across different LLMs, as well as differences in GPU resources available on various edge devices or even on the same device under different conditions, the offline component must determine the most suitable quantization strategy based on these factors. During the online phase, the selected quantization strategy is applied to quantize the LLM to the corresponding bit widths. The quantized LLM is then deployed onto the target device, ensuring that the LLM can operate efficiently in resource-limited environments while maintaining a high level of accuracy.
@@ III-A Layer Importance Detection Module
The Layer Importance Detection (LID) module identifies the key layers that have the greatest impact on LLM performance by analyzing the structure and behavior of the LLM. This ensures that essential functions of the LLM are preserved during quantization.
To assess the layer importance, we utilize intrinsic model features as metrics. We analyze each layer of the LLMs, extracting the hidden states of the last token, denoted as $X_{i,in}$ and $X_{i,out}$. These hidden states capture the encoded information for the last token in the input sequence at that particular layer. Next, we multiply these hidden states by the transpose of the matrix $W_{E}$. The matrix $W_{E}$ serves as an input embedding matrix, which is capable of mapping high-dimensional input data into a lower-dimensional vector space, thereby extracting the key features of the hidden states. We can directly obtain the matrix from the corresponding LLM. This multiplication operation produces a projection of each token across the vocabulary.
Following this, we sort the results of the projection in descending order. By selecting the top-$k$ indices from the sorted results, we identified the top-$k$ most probable tokens for the input and output of the current layer. Among them, the selection of the $k$ value is based on the work of Li et al [25]. However, the specific value chosen for $k$ has minimal impact on the overall trend. This is because, in reality, smaller $k$ values are essentially subsets of larger $k$ values.
Finally, we convert these indices back into actual tokens, forming top-$k$ token sets that represent the input and output of the layer. Figure 2 illustrates the detailed process of constructing the top-$k$ token sets.
@@
After obtaining the two sets $C_{i,in}$ and $C_{i,out}$ through Equ.(1) and Equ.(2), we calculate the Jaccard similarity [26] between the two sets as a metric for assessing the importance of the layer. Jaccard similarity has been widely used in various fields (e.g., machine learning and text mining) to evaluate the similarity between sets of samples or features. This similarity is calculated based on the ratio of the size of the intersection of the above two sets to the size of their union. This metric is simple yet effective, providing a clear measure of the compositional similarity between the two top-$k$ token sets.
@@
According to Equ.(3), we employ the Jaccard similarity to quantify the semantic transformation degree of each layer in LLMs. Specifically, we calculate the Jaccard similarity between two top-$k$ token sets derived from the input and output of each layer. Then, we transform this value by inversion and addition to obtain the importance metric $I_{i}$ for each layer. As the Jaccard similarity increases, the importance metric $I_{i}$ decreases correspondingly.
Figure 3: Importance of LLMs layer.
This observation indicates that a higher similarity between the two top-$k$ token sets implies that the layer may have performed less significant semantic transformation on the input information. Consequently, it can be inferred that the layer’s contribution to semantic transformation is relatively low, leading to a lower assessment of its importance. Through Figure 3, we can visually observe the contribution of different layers to the overall performance of the Llama-2-7B/13B [7] and Llama-3-8B [27] within this evaluation framework.
Compared to directly using cosine similarity, our approach offers a more fine-grained analysis by directly mapping hidden states to the vocabulary. This method captures the relationship between each layer and semantic information more effectively, helping to reveal the features and aspects that each layer focuses on. As a result, it enables a more accurate assessment of the importance of the layer.